OpenStudy (animelover):

Each club member collected 15 boxes of recycling for community service. Let C represent the number of club members and T represent to total number of boxes collected. Which equation models the situation? T + 15 = C 15C = T 15T = C C15=T

OpenStudy (sstradic):

second answer choice ;)

OpenStudy (animelover):

Thank you!

OpenStudy (sstradic):

your so very welcome! ;D

OpenStudy (animelover):

Can you help me with another?

OpenStudy (animelover):

Which equation represents the relationship shown in the table? x 1 5 7 y 4 20 28 x=y−15x=y−15 y=x4y=x4 y=x+3y=x+3 y=4x